# On five questions of Bourin and Lee: symmetric moduli and an Euler operator identity

### v2

@@ 1.1. A problem on matrix weighted parallelogram law
We first answer Question 1.2 in the negative: the restriction $0\leq x\leq 1$ in Theorem 1.1 cannot be removed and therefore is sharp (see Theorem 1.3 below).
@@
1.2. Symmetric moduli and three related questions
For two Hermitian matrices $A,B\in\mathbb{M}_{n}$, we write $A\leq B$ if BAB-A is positive semidefinite. While for $A,B\in\mathbb{M}_{n}$ one has the triangle inequality
$\|A+B\|_{1}\leq\|A\|_{1}+\|B\|_{1},$
simple examples show that the inequality
$|A+B|\leq|A|+|B|$
need not hold in the positive semidefinite order. However, Thompson [20] observed that there exist unitaries $U,V\in\mathbb{M}_{n}$ such that
$|A+B|\leq U|A|U^{*}+V|B|V^{*},$ (1.1)
which, upon taking traces, yields the triangle inequality as a consequence. Such “unitary-orbit” inequalities have since been generalized and extended to a variety of settings; see the surveys by Bourin, Harada and Lee [7, 9].
In this paper, we generalize (1.1) to the setting of rectangular matrices (see Lemma 5.2) in order to address Question 1.11.
Recall the Cartesian decomposition of a matrix $Z\in\mathbb{M}_{n}$:
$Z=\operatorname{Re}Z+i\,\operatorname{Im}Z,\qquad\operatorname{Re}Z:=\frac{Z+Z^{*}}{2},\quad\operatorname{Im}Z:=\frac{Z-Z^{*}}{2i}.$
Note that $|\operatorname{Re}Z|^{2}+|\operatorname{Im}Z|^{2}=\left(|Z|^{2}+|Z^{*}|^{2}\right)/2$. Bourin and Lee [12, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus, which can be regarded as a matrix analogue of the scalar inequality $|z|\leq|a|+|b|$ for a complex number $z=a+ib$.
@@ 1.2. Symmetric moduli and three related questions
They also posed the following question concerning the optimality of the exponent $2$.
Question 1.5 ([12, Question 4.6]).
@@
Our second result answers Question 1.5 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
@@
For $n=1$, the statement of Question 1.5 is different: for a complex number $z$, the left-hand side of (1.2) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$. Hence (1.2) holds for all $p>0$ in dimension 11. The necessity $p\leq 2$ is therefore a genuinely noncommutative phenomenon requiring $n\geq 2$.
@@
In [12, Corollary 4.7], the indices on the right-hand side appear with a typographical slip. Starting from the operator inequality in [12, Corollary 4.3] and applying the Weyl inequality for singular values,
@@
one obtains the corrected bound
$\mu_{1+j+k}^{\downarrow}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)\leq\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z).$ (1.3)
Accordingly, the natural corrected version of [12, Question 4.8] asks whether there exists $Z\in\mathbb{K}$ such that for every $p>2$ one has
@@
We answer Question 1.9 in the affirmative.
@@
Let
$|Z|_{\mathrm{sym}}:=\frac{|Z|+|Z^{*}|}{2},\qquad|Z|_{\mathrm{qsym}}:=\left(\frac{|Z|^{2}+|Z^{*}|^{2}}{2}\right)^{1/2}$
be the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$, respectively.
Now, we turn to a related problem raised by Bourin and Lee concerning Thompson-type triangle inequalities for symmetric moduli, compared with (1.1).
Question 1.11 ([12, Question 4.9]).
Denote by $|Z|_{\mathrm{(q)sym}}$ either the arithmetic symmetric modulus or the quadratic symmetric modulus of $Z\in\mathbb{M}_{n}$. Let $X,Y\in\mathbb{M}_{n}$. Does the Thompson triangle type inequality
$|X+Y|_{\mathrm{(q)sym}}\leq U\,|X|_{\mathrm{(q)sym}}\,U^{*}+V\,|Y|_{\mathrm{(q)sym}}\,V^{*}.$ (1.5)
hold for some unitary $U,V\in\mathbb{M}_{n}$ ?
Our next theorem provides a complete answer to Question 1.11: the inequality (1.5) fails for the arithmetic symmetric modulus but holds for the quadratic symmetric modulus.
Theorem 1.12.
Let $X,Y\in\mathbb{M}_{n}$.
  1. (i)
    In general, there do not exist unitaries $U,V$ such that
    $|X+Y|_{\mathrm{sym}}\leq U|X|_{\mathrm{sym}}U^{*}+V|Y|_{\mathrm{sym}}V^{*}.$
  2. (ii)
    There exist unitaries $U,V$ such that
    $|X+Y|_{\mathrm{qsym}}\leq U|X|_{\mathrm{qsym}}U^{*}+V|Y|_{\mathrm{qsym}}V^{*}.$
1.3. Matrix analogues of Euler’s quadrilateral identity
Finally, Bourin and Lee [12] asked for matrix analogues of Euler’s quadrilateral identity.
Question 1.13 ([12, Question 4.10]).
@@ 1.3. Matrix analogues of Euler’s quadrilateral identity
We recall that an isometry is a rectangular matrix $W\in\mathbb{M}_{m,n}$ such that $W^{*}W=I_{n}$. Bourin and Lee [12] suggested exploring isometry/unitary-orbit refinements of Euler-type identities. Our following three results give isometry-orbit refinements of (1.6).
@@
Unfortunately, the answer to Question 1.18 is negative. In fact, take $A=B=C=I_{n}$. Then

### v3

# An operator triangle inequality for the quadratic symmetric modulus

@@ 1.4. Questions of Bourin and Lee on symmetric moduli and Euler-type identities
We first answer Question 1.26 in the negative: the restriction $0\leq x\leq 1$ in Theorem 1.25 cannot be removed and therefore is sharp (see Theorem 1.27 below).
@@
Bourin and Lee [10, Corollary 4.3] obtained the following unitary-orbit estimate for the quadratic symmetric modulus, which can be regarded as a matrix analogue of the scalar inequality $|z|\leq|a|+|b|$ for a complex number $z=a+ib$.
@@
Clearly, since the function $t\mapsto t^{p/2}$ is operator concave on $0,\infty)$ for $0<p<2$, we have
$\frac{|Z|^{p}+|Z^{*}|^{p}}{2}=\frac{(|Z|^{2})^{p/2}+(|Z^{*}|^{2})^{p/2}}{2}\leq\left(\frac{|Z|^{2}+|Z^{*}|^{2}}{2}\right)^{p/2}.$
Since both sides are positive semidefinite and the function $t\mapsto t^{1/p}$ is increasing on $0,\infty)$, Weyl’s monotonicity theorem yields
$\lambda_{j}\!\left(\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\right)\leq\lambda_{j}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)$
for all $j$, where $\lambda_{j}(\cdot)$ denotes the $j$-th largest eigenvalue. Therefore,
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq W\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\,W^{*}$
for some unitary matrix $W$. Therefore, by Theorem 1.28, for any $0<p<2$, we obtain
$\left(\frac{|Z|^{p}+|Z^{*}|^{p}}{2}\right)^{1/p}\leq U|\operatorname{Re}Z|U^{*}+V|\operatorname{Im}Z|V^{*}.$
Bourin and Lee [10] then raised the following question concerning whether the exponent $2$ is optimal.
Question 1.29 ([10, Question 4.6]).
@@
Our next result answers Question 1.29 affirmatively in every noncommutative dimension $n\geq 2$ by constructing a $2\times 2$ counterexample for $p>2$.
@@
For $n=1$, the statement of Question 1.29 is different: for a complex number $z$, the left-hand side of (1.10) equals $|z|$ for every $p>0$, while the right-hand side equals $|\Re z|+|\Im z|\geq|z|$. Hence (1.10) holds for all $p>0$ in dimension 11. The necessity $p\leq 2$ is therefore a genuinely noncommutative phenomenon requiring $n\geq 2$.
@@
In [10, Corollary 4.7], the indices on the right-hand side appear with a typographical slip. Starting from the operator inequality in [10, Corollary 4.3] and applying the Weyl’s inequality for singular values,
@@
where $\mu_{j}^{\downarrow}(\cdot)$ denotes the $j$-th largest singular value, one obtains the corrected bound
$\mu_{1+j+k}^{\downarrow}\!\left(\sqrt{\frac{|Z|^{2}+|Z^{*}|^{2}}{2}}\right)\leq\mu_{1+j}^{\downarrow}(\operatorname{Re}Z)+\mu_{1+k}^{\downarrow}(\operatorname{Im}Z).$
Accordingly, the natural corrected version of [10, Question 4.8] asks whether there exists $Z\in\mathbb{K}$ such that for every $p>2$ one has
@@
We answer Question 1.33 in the affirmative.
@@
Bourin and Lee [10] finally asked for matrix analogues of Euler’s quadrilateral identity.
Question 1.35 ([10, Question 4.10]).
@@
Bourin and Lee [10] suggested exploring isometry/unitary-orbit refinements of Euler-type identities. Our following three results give isometry-orbit refinements of (1.12).
@@
Remark 1.40.
The Euler-type orbit identities in Proposition 1.36 and Theorems 1.38 and 1.39 also admit quadratic symmetric modulus versions. Indeed, these two results extend verbatim to rectangular matrices with a fixed number of columns, since their proofs only involve the positive matrices $|A|^{2}=A^{*}A$ and the Euler identity (1.12), which remains valid in the rectangular setting. Applying these rectangular versions to the lifting $\Gamma(T):=\frac{1}{\sqrt{2}}\binom{T}{T^{*}}\in\mathbb{M}_{2n,n},$ we obtain the corresponding analogues for quadratic symmetric modulus.
@@
Unfortunately, the answer to Question 1.41 is negative. In fact, take $A=B=C=I_{n}$. Then